### v1

# QPanda3: A High-Performance Software-Hardware Collaborative Framework for Large-Scale Quantum-Classical Computing Integration

Abstract
QPanda3 is a high-performance quantum programming framework that enhances quantum computing efficiency through optimized circuit compilation, an advanced instruction stream format (OriginBIS), and hardware-aware execution strategies. These engineering optimizations significantly improve both processing speed and system performance, addressing key challenges in the NISQ era. A core innovation, OriginBIS, accelerates encoding speeds by up to 86.9× compared to OpenQASM 2.0, while decoding is 35.6× faster, leading to more efficient data handling, reduced memory overhead, and improved communication efficiency. This directly enhances the execution of quantum circuits, making large-scale quantum simulations more feasible. Comprehensive benchmarking demonstrates QPanda3’s superior performance: quantum circuit construction is 20.7× faster, execution speeds improve by 3.4×, and transpilation efficiency increases by 14.97× over Qiskit. Notably, in compiling a 118-qubit W-state circuit on a 2D-grid topology, QPanda3 achieves an unprecedented 869.9× speedup, underscoring its ability to handle complex quantum workloads at scale. By combining high-speed quantum processing with a modular and extensible software architecture, QPanda3 provides a practical bridge between today’s NISQ devices and future fault-tolerant quantum computing. It facilitates real-world applications in financial modeling, materials science, and combinatorial optimization, while its robust and scalable design supports industrial adoption and cloud-based deployment.
Index Terms:
Quantum Computing, Software-Hardware Collaborative, High-Performance, Quantum Circuit Compilation, Intermediate Representation
Figure 1: Comparison of average time utilized for various quantum circuit processing tasks between QPanda3 and Qiskit. The tasks include circuit compilation on different topologies (all-to-all, square, heavy-hex, and linear), circuit manipulation, and circuit construction. The results indicate that QPanda3 demonstrates significantly lower compilation time across all topologies compared to Qiskit, while the time difference in manipulation and construction tasks is relatively smaller.
I Introduction
Quantum computing, as a revolutionary computational paradigm, has garnered widespread attention from both academia and industry due to its exceptional parallel computing capabilities and potential to solve problems intractable for classical computers. In recent years, significant advancements in quantum hardware have propelled quantum computing from theoretical exploration to practical applications. However, to fully harness its potential, robust software frameworks are essential to bridge the gap between users and quantum hardware. The effectiveness of quantum computing software frameworks depends on several key factors, including user accessibility, computational efficiency, and seamless hardware integration. Against this backdrop, QPanda3 [1], a modern quantum programming framework, has undergone extensive optimization to enhance usability, quantum circuit compilation efficiency, and program transmission performance, thereby driving quantum computing toward broader applications.
Currently, quantum computing is undergoing a transition from theoretical research to engineering practice. Despite the challenges faced by existing quantum hardware, such as noise, decoherence, and limited qubit scalability, engineering optimizations are steadily improving its usability, enabling Noisy Intermediate-Scale Quantum (NISQ) devices to demonstrate computational advantages over classical computers for specific problems. This engineering evolution not only determines the practical applications of NISQ devices but also influences the feasibility of large-scale fault-tolerant quantum computing in the future. Thus, enhancing the usability, stability, and execution efficiency of quantum computing has become a central challenge in its research and application. Key approaches to addressing these challenges include compilation optimization, instruction stream design, and software-hardware co-optimization.
As a high-performance quantum programming framework, QPanda3 adopts an engineering-driven approach to enhance the practical usability of quantum computing. Its core objectives include:
  • Enhancing the computational capabilities of existing quantum devices—Through efficient compilation techniques, optimized qubit mapping strategies, and quantum gate compression algorithms, QPanda3 enables NISQ devices to execute target algorithms more efficiently, achieving superior computational performance for specific tasks.
  • Optimizing software-hardware co-design in quantum computing—Unlike classical computing, quantum computing requires software to have deep awareness of hardware characteristics. QPanda3 employs adaptive qubit mapping and hardware-aware compilation strategies to ensure optimal execution across different quantum processors, thereby improving overall computational fidelity.
  • Advancing quantum computational advantages—Although some NISQ devices have demonstrated computational superiority for certain problems, achieving large-scale quantum computing necessitates further engineering breakthroughs. QPanda3 adopts a modular architecture, allowing seamless integration with future fault-tolerant quantum computing systems, thereby reducing the transition cost from NISQ to next-generation quantum technologies.
The intermediate representation (IR) of quantum programs serves as a crucial bridge between high-level quantum programming languages and low-level hardware instructions, influencing program readability, portability, and execution efficiency. However, traditional quantum IR solutions often suffer from poor readability and low transmission efficiency, making it difficult for researchers to understand the mapping between high-level algorithms and low-level hardware execution. Furthermore, the lack of a unified IR standard among mainstream quantum programming languages hinders cross-platform algorithm benchmarking and experimental comparisons.
To address these issues, QPanda3 introduces a novel IR framework featuring OriginIR and an efficient binary instruction set OriginBIS to enhance program readability, transmission efficiency, and cross-platform compatibility. In classical-quantum hybrid computing and multi-device collaborative computing scenarios, existing IR solutions are typically optimized for specific hardware architectures, limiting their portability across different quantum platforms. By optimizing its IR design, QPanda3 enables efficient migration across diverse quantum platforms, thereby improving task allocation flexibility and execution efficiency.
Efficient quantum circuit compilation is crucial for transforming abstract quantum algorithms into hardware-executable instructions, encompassing quantum gate decomposition, qubit mapping, and circuit optimization. Building upon its predecessor, QPanda2, QPanda3 introduces multiple optimizations that significantly enhance compilation efficiency. Benchmarking results indicate that QPanda3 surpasses QPanda2 and other mainstream quantum computing frameworks in circuit compilation speed, dramatically improving quantum algorithm execution efficiency. In the NISQ era, where hardware constraints are prevalent, optimizing compilation efficiency is critical for maximizing computational utility.
As a quantum programming framework focused on software-hardware co-optimization, QPanda3 excels in cloud-based quantum computing environments. Given the scarcity and high cost of quantum computing resources, centralized cloud deployment is often adopted to enhance accessibility and resource utilization. Additionally, cloud-based quantum computing facilitates the establishment of standardized software-hardware environments, allowing users to conduct experiments and development without direct access to quantum hardware. Moreover, the cloud computing model fosters ecosystem collaboration in quantum computing, simplifying resource sharing and algorithm optimization.
To achieve optimal quantum execution, a comprehensive understanding of quantum hardware characteristics—such as qubit connectivity, error rates, and coherence times—is essential. QPanda3 integrates device-aware compilation techniques to minimize circuit depth and gate operations, thereby improving computational efficiency. Furthermore, QPanda3 provides device-specific quantum program analysis tools to help users assess execution performance, identify potential bottlenecks, and further optimize quantum programs. Through software-hardware co-optimization, QPanda3 enables quantum programs to adapt to different hardware architectures, achieving more efficient computation.
Experimental results demonstrate that QPanda3 significantly outperforms existing quantum programming frameworks across multiple performance metrics. For example:
  • Compilation efficiency: Experiments on Benchpress[2], in a fully connected topology, QPanda3’s compilation speed is, on average, 9.46× faster than Qiskit[3], with peak acceleration reaching 77.76×. In a square topology, its average compilation speed surpasses Qiskit[3] by 24.6×, with a peak of 869.9×. For heavy-hexagon topologies, QPanda3 achieves an average 15.1× speedup, with a peak of 332.0×. In linear topology, the average speedup is 10.71×, with peak acceleration reaching 868.7×.
  • IR transmission efficiency: The encoding speed of OriginBIS is 18.7× faster than OriginIR and 86.9× faster than OpenQASM 2.0 [4]. The decoding speed is 18.9× faster than OriginIR and 35.6× faster than OpenQASM 2.0.
The content of this paper is organized as follows: Firstly, we review the related work on quantum computing compilation frameworks. Secondly, we provide an overview of the overall architecture and core features of QPanda3. Then, we successively introduce the representation method and transmission mechanism of quantum programs, device-based quantum circuit optimization and compilation strategies, and device-based quantum program analysis, delving into the three key components of QPanda3. Finally, we validate the performance advantages of QPanda3 through experiments.
II Related Work
II-A Intermediate Representation of Quantum Programs
The Intermediate Representation (IR) of quantum programs serves as a crucial bridge connecting high-level quantum algorithm descriptions with underlying hardware implementations in quantum computing. IR provides descriptions of quantum programs at various levels of abstraction, each suited for the design, processing, transmission, and execution of quantum programs. Lower-level abstractions tend to be closer to the machine instruction format of quantum computing devices, while higher-level abstractions are more easily understood by humans. Representing quantum programs based on the quantum circuit model is a common approach, and most IRs of quantum programs are built upon this foundation. Even for IRs of quantum programs that include classical instructions and complex control flows, they are typically extensions based on the quantum circuit model.
Quantum Assembly Language (QASM)[5] is an important form of intermediate representation for quantum programs, primarily used to provide a low-level abstraction of quantum programs. It is important to note that QASM for quantum computing differs significantly from assembly languages in classical computer architectures. Although QASM is a low-level representation, it still requires the device to dynamically manage quantum registers and classical registers (logical-to-physical mapping) rather than directly operating on hardware registers.
Intermediate representations of quantum programs that resemble high-level programming languages offer good user readability. However, such representations tend to be relatively complex. For example, with the increasing complexity of quantum computing systems and application demands, IBM has continuously extended QASM to develop OpenQASM 2.0 and OpenQASM 3.0. As widely adopted versions, OpenQASM 2.0[4] and OpenQASM 3.0[6] not only extend quantum gate operations and qubit management but also introduce more complex syntactic structures and functional features, such as classical control flow and modular programming support. They are more akin to the style of the classical programming language C compared to the original QASM. F-QASM[7] extends QASM with feedback instructions, enhancing the efficiency of implementing measurement-based branch and loop statements. Similar intermediate representations include Scaffold[8], QCL[9], Quipper[10], ProjectQ[11], and QIR[12]. These types of intermediate representations for quantum programs sacrifice portability due to their complexity, hindering comparative experimental research by researchers.
Device-oriented intermediate representations of quantum programs exhibit strong dependencies on the specific device. OpenPulse[13], Pulser[14], and JaqalPaw[15] are examples of intermediate representations tailored for pulse-based devices, offering fine-grained control over quantum devices. QuMIs[16] is an intermediate representation designed for distributed device control. eQASM[17], on the other hand, is a low-level intermediate representation that is directly linked to binary machine instructions. Clearly, these intermediate representations contain a significant amount of device-specific information within the code describing quantum programs, which provides precise control over the execution details of quantum programs but also reduces their device independence.
The intermediate representation of quantum programs can also serve as a transmission protocol for transferring quantum programs between devices. Currently, there is relatively little research discussing this topic. It is important to note that this type of protocol differs from NetQASM[18] and InQuIR[19]. NetQASM and InQuIR are quantum network-oriented intermediate representations based on quantum communication protocols, used to control the devices and processes involved in quantum communication.
II-B Compilation, Optimization, Qubit Mapping, and Routing of Quantum Circuits
Quantum logic circuits can represent the quantum computing component of quantum programs. The quantum circuit model can also depict the actual physical operation steps in a quantum processor. The quantum software stack provides high-level programming languages for designing the former. The latter typically corresponds to a sequence of machine instructions that directly operate the quantum device. Quantum circuit compilation is the process of converting quantum logic circuits into sequences of machine instructions. The transformation from an initial quantum logic circuit to a sequence of machine instructions involves multiple structural conversions of the quantum circuit. These conversions include steps such as unitary matrix decomposition, qubit mapping, qubit routing, optimization, and compilation into a sequence of machine instructions.
Quantum computing software such as Qiskit and QASMTrans[20] provides modern support for quantum circuit compilation. Paulihedral[21] is a compiler designed specifically for the VQE algorithm. Similarly, application-oriented compilers also include Twoqan[22], which is dedicated to QAOA circuits. CaQR focuses on the generation of dynamic circuits[23], is a compiler specifically for pulse-based devices. AutoComm[24] and QuComm[25], on the other hand, provide support for distributed devices.
Qubit mapping and routing are essential steps for quantum logic circuits to be executed by quantum processors. Qubit mapping assigns corresponding physical qubit resources to each logical qubit, while qubit routing ensures the connectivity constraints between physical qubits. Optimization aims to obtain compiled circuits with good computational performance. Sabre[26] proposes a method to minimize the number of ancillary qubits. TOQM[27], on the other hand, aims to reduce the depth of the translated circuits[28], among others, investigates optimization schemes related to quantum gate aggregation.
II-C Quantum Program Profiling
Quantum program profiling is of great significance for fully harnessing the potential of quantum computing. With the development of quantum software, quantum programs often contain not only pure quantum gate operations but also other classical instruction execution processes. Therefore, quantum program profiling involves the analysis of quantum circuits and operational subprocesses.
Quantum computing performance analysis benchmarks can be utilized for quantum circuit analysis. Currently, numerous proposed benchmarks are employed for quantum computing performance analysis, primarily focusing on evaluating the performance of quantum processors when executing specific circuits. These metrics include quantum volume[29], Q-score[30], quantum LINPACK[31], and quantum process tomography[32], as well as using quantum applications like VQE[33] and QAOA[34] to analyze the performance of quantum processors. SupermarQ further proposes multiple analysis benchmarks. Clearly, these benchmarks also reflect the performance of the quantum circuits used for testing on the execution device. The benchmarks proposed by SupermarQ[35], such as inter-qubit communication volume, critical depth, coherence rate, parallelism, and qubit activity, can be calculated based on available device information and quantum circuit structure, and thus can also be applied to device-performance-oriented quantum circuit analysis.
Ideas from classical program analysis can be applied to analyze the operational processes in quantum programs. Qprof[36], based on gprof[37], can be used for quantum program analysis to obtain metrics such as subprocess invocation rates, qubit occupancy rates, and time consumption. In this paper, QPanda3 is introduced, which applies process analysis methods to the interconnectivity analysis of quantum logic gates and qubit utilization analysis.
III Overview of QPanda3
QPanda3 is an advanced quantum programming framework that provides comprehensive support for users in quantum computing across both software and hardware aspects. As illustrated in Figure 2, QPanda3 not only offers abstractions for quantum programs and computing devices but also provides multiple related components and tools.
III-A Oriented Towards Quantum Computing
The core of quantum computing software lies in quantum programs, for which QPanda3 provides multi-level abstractions. Specifically, QPanda3 describes quantum programs using a quantum gate-based circuit model, while distinguishing between circuits without measurements and those with measurements to emphasize the difference between measurement operations and quantum gate operations. This distinction is directly reflected in QProg and QCircuit, which are part of the high-level programming language mechanism provided by QPanda3. The high-level programming language mechanism of QPanda3 uses instruction sequences within Python scripts to represent quantum programs. Users can design such quantum programs using Python and the APIs provided by QPanda3. Although high-level programming languages like Python are more user-friendly, they also present issues such as complex mapping relationships with hardware machine instructions and low transmission efficiency. QPanda3 offers a streamlined and efficiently transmittable intermediate representation of quantum programs. Furthermore, QPanda3 can compile these programs into a machine-level instruction stream format for quantum programs.
QPanda3 abstracts quantum computing hardware to meet users’ needs for executing quantum programs and conducting hardware-software co-design. In terms of computing resources, QPanda3 abstracts real quantum processors, quantum simulators, and classical processors, thereby facilitating users to efficiently utilize these resources to complete computational tasks. Regarding the physical attributes of quantum systems, QPanda3 provides noise simulation and Hamiltonian simulation, supporting users in conducting related experimental research.
Centered around quantum programs, QPanda3 offers a rich set of components. A significant portion of these components is designed to meet the needs of hardware-software co-design, including modules specifically for quantum circuit optimization and compilation, modules tailored for device-based quantum program analysis, and application interfaces that facilitate local-cloud collaboration. Other components provide various supports for quantum programs, primarily including a quantum program translation module to ensure program portability, a quantum program visualization module to enhance design efficiency, and a variational quantum circuit module for batch generation of circuits and post-processing.
Figure 2: Framework
QPanda3 aims to fully satisfy users’ requirements for quantum computing-related work and therefore provides numerous useful tools. These tools pertain to quantum information, operators, and testing convenience. The tools related to quantum information offer multiple representations of quantum states and quantum channels, along with several information analysis utilities. The tools concerning operators provide corresponding transformation and simulation tools. The testing tools include interfaces for randomly generating quantum circuits, which can produce a large number of random circuits based on different configurations, thereby meeting users’ diverse testing needs.
III-B From QPanda2
QPanda3 fully leverages the successful experience of QPanda2 in terms of design and functionality. It continues the support for the Python language, enabling researchers and developers to quickly implement quantum algorithms using Python’s concise syntax and rich ecosystem. This design not only lowers the threshold for quantum programming but also enhances development efficiency, making it particularly suitable for beginners in quantum computing and cross-disciplinary researchers. QPanda3 introduces OriginIR as an efficient intermediate representation for quantum programs and provides conversion tools between OriginIR and OpenQASM. This design facilitates seamless migration of quantum programs across different frameworks and hardware platforms, while also providing convenience for the optimization and analysis of quantum circuits. QPanda3 supports multiple types of quantum simulators (such as full amplitude simulators and partial amplitude simulators), quantum noise simulation, Hamiltonian simulation, and management of classical registers. This comprehensive hardware abstraction capability allows researchers to verify and optimize quantum algorithms in various scenarios. To reduce the learning curve for users, QPanda3 retains the quantum circuit construction methods consistent with QPanda2. Users can quickly get started with familiar interfaces and syntax, while enjoying the performance improvements and functional extensions offered by the new framework.
III-C Enhanced Performance
QPanda3 demonstrates significant performance improvements, with notable advantages primarily manifesting in efficient quantum program transmission and high-performance quantum circuit compilation.
(1)Efficient Quantum Program Transmission
QPanda3 has designed a set of efficient protocols specifically for quantum program transmission, aiming to optimize the process of transmitting quantum programs between terminal and cloud devices, as well as among other devices. By standardizing the representation format and transmission process of quantum programs, this protocol significantly enhances the efficiency of users submitting tasks to quantum computing devices. Specifically, the protocol supports encapsulating and transmitting quantum programs in a compact intermediate representation form, avoiding the inefficiencies caused by non-uniform formats or excessive redundant information in traditional transmission methods. Additionally, the protocol employs efficient compression and encoding techniques, effectively reducing the amount of data during transmission, thereby saving communication bandwidth between devices and accelerating interaction speed. In practical applications, this protocol is not only suitable for task submission between user terminals and cloud-based quantum computing devices but also for program sharing and collaborative computing among different quantum computing devices. For example, in distributed quantum computing scenarios, multiple quantum devices can quickly exchange quantum programs and data through this protocol, achieving more efficient parallel computing and resource utilization. Furthermore, the design of this protocol takes into account the compatibility and scalability of quantum programs, adapting to the evolution needs of future quantum hardware and algorithms.
(2) Efficient Quantum Circuit Optimization and Compilation
The implementation and optimization of quantum algorithms still face numerous challenges, particularly in the construction and compilation of quantum circuits. As an advanced quantum computing framework, QPanda3 is dedicated to addressing these issues and providing efficient quantum circuit compilation tools. Through optimized implementation, QPanda3 significantly accelerates the process of constructing quantum logic circuits from Python code, enhancing the development efficiency of quantum algorithms. Secondly, the framework introduces the Transpiler lass, which enables device-oriented quantum circuit compilation. It dynamically adjusts compilation strategies based on the characteristics of the target hardware, thereby improving the execution efficiency and portability of quantum programs. Furthermore, QPanda3 optimizes the quantum circuit compilation process from multiple aspects, significantly enhancing the efficiency of quantum circuit compilation. With efficient quantum circuit compilation technology, QPanda3 provides strong support for the implementation and optimization of quantum algorithms.
III-D Software-Hardware Co-Design Oriented
QPanda3 is a modern quantum computing software development kit oriented towards software-hardware co-design. This is manifested in various key scenarios including quantum program design, quantum program transmission, quantum circuit compilation, quantum program analysis, and quantum program execution. As illustrated in Figure 3, QPanda3, running on both terminal and cloud devices, fully meets the critical requirements of quantum computing based on software-hardware co-design.
Figure 3: Workflow
(1) Cloud Execution and Inter-device Transmission of Quantum Programs
Due to the unique characteristics of quantum computers, the paradigm of terminal design for quantum programming coupled with cloud execution has been widely accepted and has found numerous practical applications. A hallmark of this paradigm is the separation of devices used for program design from those used for program execution, with the two interconnected and collaborating through communication links. Quantum programs are transmitted from the design end to the computation end via these communication links. QPanda3 is not merely a tool for quantum programming; rather, it is a collaborative tool that adheres to the aforementioned paradigm. On one hand, QPanda3 provides convenient access to the quantum cloud, allowing quantum programs designed on QPanda3 to be directly executed on various computing devices within the quantum cloud. On the other hand, to optimize the efficiency of quantum program transmission between devices, QPanda3 has designed a compact intermediate representation for quantum programs. Additionally, QPanda3 offers tools for converting between quantum programs expressed in high-level programming languages and this intermediate representation. Furthermore, serving as an interactive interface for users to access quantum computing devices, QPanda3 enables users to obtain relevant information about the computing devices.
(2) Hardware-Software Co-Design for Optimized Compilation of Quantum Circuits
Quantum circuit compilation is the process of transforming quantum logic circuits into sequences of machine instructions. This process, which is a software-based procedure, necessitates a comprehensive consideration of the actual connection topology of physical qubits and the physical operations that the device employs to control these qubits, with such operations typically corresponding to a limited set of quantum gates. Clearly, quantum circuit compilation represents a typical scenario of hardware-software co-design. The extent of various optimizations applied during the quantum circuit compilation process influences the degree to which quantum programs utilize device performance, and directly impacts the correctness of quantum computation results. QPanda3 decouples key steps in quantum circuit compilation, including qubit mapping, qubit routing, and machine instruction generation, thereby enabling users to compile and generate machine instruction sequences that can be efficiently executed on specified devices, based on different physical qubit layouts and native quantum gate sets. During the circuit compilation process, QPanda3 implements multiple software-level optimization strategies, some of which accelerate compilation speed, while others enhance the execution performance of the compiled circuits.
(3) Hardware-Software Co-Design for Quantum Program Analysis and Design
The goal of quantum programming is to develop quantum programs that can solve target problems, fully utilize computational resources, and exhibit acceptable quantum computing performance. For the quantum computing component, the key to quantum programming lies in obtaining efficient quantum logic circuits.
QPanda3 enables users to design quantum programs using a high-level programming language (Python) and a user-friendly intermediate representation (OriginIR). QPanda3 provides a variety of quantum simulators, which can be deployed on the quantum cloud as services for users. These simulators are also built into QPanda3, allowing users to run them using the computing resources of their personal computers (CPUs, GPUs). Quantum simulators on the cloud and on personal computers can be used to verify whether the theoretical execution results of quantum programs match expectations, thereby guiding the further design of quantum programs. Collaborating with local or cloud computing resources to validate the correctness of quantum programs represents a classic hardware-software co-design scenario. There may be multiple implementations of quantum programs that solve the same problem, and different quantum programs may exhibit varying performance on the same device. It is of great significance to select an appropriate implementation (or subroutine) of a quantum program that is tailored to the target problem and the available computing device. Such task- and device-adapted quantum programs can obtain more accurate results more efficiently, aligning more closely with the goals of quantum programming. This type of quantum program design relies on device-based quantum program analysis. QPanda3 provides this analysis capability, facilitating users in designing more efficient quantum programs. Specifically, QPanda3 offers quantum program analysis tools that are oriented towards quantum computing performance.
IV Representation and Transmission of Quantum Programs
IV-A User-Friendly Intermediate Representation – OriginIR
The significant role of high-level intermediate representations of quantum programs in the quantum software stack cannot be overstated. QPanda3 employs OriginIR as a lower-level representation of quantum programs compared to high-level programming languages.
IV-A1 Compatibility with QPanda2
QPanda3 largely retains all the features of OriginIR from QPanda2, while also introducing appropriate extensions. Specifically, the syntax for declaring classical bits, declaring quantum bits, separating adjacent instruction statements, representing block statement scopes, and declaring quantum logic gates and quantum-related operations remains consistent with QPanda2. QPanda3 also supports all the quantum logic gates available in QPanda2. OriginIR instruction strings generated by QPanda2 can be correctly parsed and processed by QPanda3. Conversely, OriginIR instruction strings generated by QPanda3, which do not include auxiliary information such as comments, can also be correctly parsed and processed by QPanda2. OriginIR enables seamless portability of quantum programs between QPanda2 and QPanda3.
IV-A2 Geared Towards Researchers
(1) Complexity and Conversion
The QASM series represents intermediate representations of quantum programs. To investigate the performance of quantum programs, researchers often test and compare the same quantum program on different platforms. QASMBench[38], based on OpenQASM-2, serves as a relevant benchmark for quantum computing-related tests and has been widely adopted in numerous research studies.
Unlike the universality of the quantum circuit model, intermediate representations (IRs) of quantum programs exhibit significant variations due to differences in quantum computing devices, abstraction levels, and syntactic structures. Even IRs of quantum programs targeting the same quantum computing device and operating at similar abstraction levels can still have considerable discrepancies. Taking the QASM series as an example, OpenQASM-2 and OpenQASM-3 differ in many syntactic aspects, to the extent that even Qiskit itself requires separate conversion tools for OpenQASM-2 and OpenQASM-3. When other software platforms provide auxiliary conversion tools, they often only support a subset of the main content within the QASM series. The more complex the syntactic rules of an intermediate representation of a quantum program are, the greater the difficulty in converting it to other quantum program IRs. This limits researchers’ ability to use quantum program IRs for porting and experimental research of quantum programs across multiple platforms. This issue has given rise to research on more universal quantum program IRs, such as cQASM[39].
OriginIR offers a perspective for reducing the difficulties associated with intermediate conversions of quantum programs, from the standpoint of low complexity. OriginIR provides a concise representation of quantum programs based on a quantum logic gate model, with minimal details unrelated to the circuit. Furthermore, OriginIR adopts a unified format for describing classical registers, as opposed to the multiple formats supported by the QASM series. Additionally, the syntactic rules of OriginIR are straightforward. These characteristics enable researchers to achieve the transplantation of quantum programs from OriginIR through simple processing.
To meet the needs of researchers, QPanda3 provides a tool for converting QASM to OriginIR. It should be noted that QPanda3 only supports a subset of the syntactic rules of OpenQASM-2. However, this tool is already capable of handling the majority of use cases in QASMBench.
(2) Readability
To enhance the readability of OriginIR, compared to QPanda2, QPanda3 allows users to freely add line comments and block comments to OriginIR instruction strings, and also permits appropriate indentation of OriginIR line instructions using spaces. Although these comments and indentations are filtered out during the process of parsing OriginIR instructions in QPanda3, the information in the comments provides important reference for users, and the indentation significantly improves the readability of the OriginIR instruction sequence. The improvement in readability will greatly facilitate researchers in conducting quantum computing-related research based on OriginIR.
IV-B Intermediate Representation for Quantum Program Transmission - OriginBIS
Quantum program transmission is a crucial component of the quantum computing software stack. Although user-friendly intermediate representations (IRs) such as OriginIR can be used to transmit quantum programs in their entirety, these IRs carry a considerable amount of redundant information. Aspects like readability are unnecessary for mere inter-machine quantum program transmission, and excessive redundant information significantly reduces the efficiency of quantum program transmission and increases unnecessary transmission costs. Compressing user-friendly IRs like OriginIR before transmission is an optional solution to reduce quantum program transmission costs. However, this approach adds compression and decompression steps, increasing the time cost of the conversion process. QPanda3 has designed a binary instruction stream (BIS), named OriginBIS, as an intermediate representation for quantum program transmission. This IR operates at the binary machine instruction level, enabling efficient quantum program transmission without adding unnecessary conversion time costs.
OriginBIS employs a stream-based approach for organization and transmission. On one hand, consecutive quantum program instructions are concatenated in data packets and transmitted over communication links in a streaming manner. On the other hand, data within individual quantum program instructions is concatenated in data packets using a streaming approach. This stream-based method allows OriginBIS to provide efficient binary representations tailored to different quantum program instructions and also supports the application of specialized optimization techniques for data transmission in quantum programs, thereby enhancing the efficiency of quantum program transmission.
OriginBIS employs a classification-based binary instruction format alignment scheme to enhance the speed at which devices process OriginBIS instructions. Specifically, OriginBIS first classifies the various instructions in a quantum program based on their function, the type and number of associated data. Then, it designs a binary instruction format with a unified length and bit fields for each category of instructions. This provides efficient support for the instruction dispatch and execution of quantum program instructions within the device. The classification strategy allows OriginBIS to provide adaptive and efficient binary instruction formats for complex instruction types, and the fact that instructions within the same category share the same format enables efficient processing of each category of instructions by the device. In this way, OriginBIS effectively balances the richness of quantum program instructions and the efficiency of device processing of these instructions.
In some quantum program transmission scenarios, the time cost of transmission may be a primary concern. To address this, OriginBIS incorporates an adaptive compression scheme. This scheme is a lightweight data compression approach based on variable-length integer encoding. It applies limited compression to the quantum program.
V Optimized Compilation of Quantum Circuits
The compilation process of quantum circuits involves the transformation from high-level abstract descriptions to hardware-executable instructions. It encompasses the entire process from initialization to final translation, with each step closely interconnected, collectively forming a complete, efficient, and reliable compilation chain. Achieving high-quality execution of quantum circuits and obtaining desirable results relies on effective adaptation and utilization of quantum computing resources. High-quality compilation of quantum circuits is a crucial approach to breaking through the performance bottlenecks in their execution. Efficient compilation of quantum circuits aims to reduce the time and economic costs associated with processing large-scale quantum circuits. The practical application of quantum software is inseparable from efficient quantum circuit compilation schemes.
V-A Transpiler of QPanda3
QPanda3 features a device-oriented modern quantum circuit transpiler. Significant differences exist among various quantum processors. Regardless of the disparities in implementation technologies, there are substantial variations in the topological structures of physical qubits among different quantum processors. Not only do different product series exhibit marked differences, but different versions within the same series may also vary due to factors such as the number of qubits. Additionally, dynamic allocation of physical qubit resources can lead to changes in the topological structure of available physical qubits on the same quantum processor. These factors pose challenges for quantum circuit transpilation. Therefore, modern quantum circuit transpilers need to generate high-quality transpiled circuits that are tailored to the current resource status of the quantum computing device. The quantum circuit transpiler built into QPanda3 takes the topological structure of available physical qubits of the quantum computing device as an input parameter and transpiles a specific quantum logic circuit into a sequence of machine instructions that can be efficiently executed on the corresponding device.
QPanda3 decouples the key steps in the internal implementation of its quantum circuit transpiler, thereby enhancing the flexibility for functional extensions of the compiler module. These key steps include preprocessing, qubit mapping, qubit routing, optimization, and machine instruction generation. The preprocessing step, as the starting point of the compilation process, is responsible for initializing the quantum circuit, including identifying the qubits, quantum gates, and their dependencies within the circuit, laying the foundation for subsequent operations. The objective of qubit mapping is to map logical qubits to the available physical qubits in the quantum processor. This process involves matching the quantum circuit with the corresponding topological structure of the physical qubits. After qubit mapping, the allocated physical qubits may not satisfy the connectivity constraints originally imposed by the quantum circuit. Qubit routing, as the immediate subsequent step to qubit mapping, is specifically designed to address this issue. Optimization is a crucial part of the quantum circuit compilation process, and reasonable optimization strategies can improve the speed and quality of compilation to varying degrees. Unlike other steps that have a strict sequential order among them, optimization may occur throughout various other steps, enhancing compilation efficiency from different aspects. Some targeted and independent optimization strategies can also be abstracted into a separate step. QPanda3 employs such a strategy, using OptimizationPass for abstraction and management, thereby facilitating multiple optimizations before qubit mapping and after qubit routing.
To promote the development of quantum computing-related research and industries, QPanda3 provides external access to its built-in transpiler through dedicated interfaces. These interfaces enable users to perform device-based quantum circuit compilation tasks. Specifically, users are required to generate the edge set of an undirected graph that corresponds to the topological structure of the available physical qubits. Apart from a few specific constraints that need to be satisfied, users can complete this step based on the information of any quantum computing device. Subsequently, users can utilize the interfaces provided by QPanda3 to generate the corresponding machine instruction sequences based on the designed quantum programs and the edge set data. Quantum programs designed using QPanda3 can seamlessly leverage these interfaces. For quantum programs developed on other platforms, users can employ the intermediate representation conversion tool provided by QPanda3 to facilitate quantum program migration. Furthermore, to further enhance user convenience, QPanda3 offers an interface for randomly generating the edge set of an undirected graph. This interface can also generate classic square, fully connected, and linear topological structures based on the number of qubits.
To achieve efficient quantum circuit compilation, QPanda3 has introduced several improvements over its predecessor, QPanda2. In addition to making the built-in transpiler accessible externally for the first time, QPanda3 has also optimized other components used in quantum program design and the quantum circuit compilation process. QPanda3 has undergone a comprehensive upgrade in its internal conversion mechanisms. During the construction of quantum circuits, extensive internal data structure and algorithm conversions are often required to ensure the correctness and executability of the circuits. QPanda3 employs more efficient data structures and algorithms, optimizing these internal conversion processes and significantly enhancing the speed of circuit construction. Whether it is the merging of quantum gates, simplification of circuits, or allocation of resources, QPanda3 can accomplish these tasks at a faster pace, thereby improving overall compilation efficiency.
V-B Starting with an Example
This section demonstrates the use of the built-in transpiler in QPanda3 through a simple example. In the following explanation, we will also introduce the basic steps for designing quantum programs using QPanda3, as well as useful tools for quantum circuit visualization and topology data generation.
(1) A Quantum Circuit Compilation Task
Figure 4, subgraph (a) illustrates the topology of available qubits on a certain quantum computing device using an undirected graph, with the corresponding physical qubits labeled as $q_{0}$, $q_{1}$, and $q_{2}$. Subgraph (b) depicts a quantum logic circuit. The compilation task in this section is to obtain the machine instruction sequence for this circuit that adapts to the given topology. Subgraph (c) presents the visualization result of the circuit in subgraph (b) using QPanda3. It is worth noting that the CZ gate is a symmetric gate.
(a) Qubits’ Topology
(b) Quantum Logic Circuit
(c) Visualization Result of The Circuit
Figure 4: Quantum circuit compilation task
QPanda3 uses ”pyqpanda3” as its Python package name, which differs from the package name ”pyqpanda” used in QPanda2. This allows users to import both the ”pyqpanda3” and ”pyqpanda” packages simultaneously, thereby ensuring compatibility with different versions. It should be noted that QPanda3 shares most of its naming conventions with QPanda2, so care should be taken to avoid issues related to name conflicts during the import process.
The core components of QPanda3 are exported by the package pyqpanda3.core. These core components include, but are not limited to, natively supported quantum logic gates such as RZ, X1, and SWAP. They also encompass the abstract object QProg for quantum programs. draw_qprog and set_print_options are two interfaces related to quantum circuit visualization tools. Specifically, set_print_options is used to control the number of decimal places displayed for gate parameters when visualizing parametric gates, while draw_qprog is used for visualizing quantum circuits.
The components related to the built-in circuit transpiler in QPanda3 are exported by pyqpanda3.transpilation. The Transpiler class is used for managing circuit compilation. The generate_topology function is used for randomly generating topology data.
from pyqpanda3.core import QProg, draw_qprog, RZ, X1, CZ, SWAP, set_print_options
from pyqpanda3.transpilation import Transpiler, generate_topology
(2) Constructing Quantum Logic Circuits
prog_1 = QProg()
prog_1 << RZ(3,1.57) << RZ(4,3.14) << RZ(5,6.28) << CZ(3,4) << CZ(4,5) << CZ(3,5)
set_print_options(2)
[…]
The output is
Figure 5: Result of Constructing Quantum Logic Circuits
(3) Setting the Topology of Physical Qubits
This line of code defines a two-dimensional Python list. This list stores the edge set of an undirected graph, which can represent the topology of the currently available physical qubits on a quantum computing device. Users can also generate similar data using the generate_topology interface.
topo = [[0,1],[1,2]]
(4) Compiling with Different Optimization Levels
The transpile method of the Transpiler object is used to perform the quantum circuit compilation step. Thanks to the generality of the quantum circuit model, a quantum circuit composed of quantum gates and other operations that correspond one-to-one with machine instructions can be mapped to an executable sequence of machine instructions. This brings many conveniences, including the ability to observe the results of quantum circuit compilation using quantum circuit visualization methods.
Due to the dynamic nature of the quantum circuit compilation process, different but equivalent compiled circuits may be generated each time. The output content is extensive and will not be shown here. In the subsequent sections, some key output results of this program will be extracted and analyzed.
transpiler = Transpiler()
prog_level_0 = transpiler.transpile(prog_1, topo, {}, 0)
[…]
V-C Qubit Mapping and Routing
(1) Qubit Mapping
Mapping logical qubits in a quantum circuit to physical qubits, taking into account the topology and connectivity of the quantum hardware, to ensure the executability of the circuit.
An example of a compilation result from the code in Section 5.2 is shown in Figure 6. Based on the parameters of the first three RZ gates added to the quantum circuit, it can be determined that logical qubit $q_{5}$ is mapped to physical qubit $q_{0}$, logical qubit $q_{4}$ is mapped to physical qubit $q_{1}$, and logical qubit $q_{3}$ is mapped to physical qubit $q_{2}$.
Figure 6: Qubit Mapping Result
(2) Qubit Routing
Inserting necessary SWAP gates between physical qubits to address connectivity limitations in quantum hardware and ensure that quantum gates can be executed smoothly on physical qubits in the predetermined order.
In the quantum logic circuit example in Section 5.2, there are direct CZ gate connections between each pair of the three logical qubits, including a direct connection between logical qubit $q_{5}$ and logical qubit $q_{3}$. However, in the aforementioned qubit mapping result, physical qubit $q_{0}$ and physical qubit $q_{1}$ are not directly connected, which violates the corresponding connectivity constraint.
In the optimization compilation at level 0 in QPanda3, the SWAP gate is implemented equivalently using the following sub-circuit.
Figure 7: Qubit Routing Result
Comparing the two figures below, it can be observed that the compilation effect is equivalent to using SWAP gates to satisfy the connectivity constraints. The equivalent circuit corresponding to the aforementioned compilation result is shown in the figure below. Clearly, this circuit is equivalent to the original quantum logic circuit. It should be noted that the example uses gates such as RZ and CZ, which are not decomposed or transformed during the preprocessing stage.
Figure 8: Using SWAP Gates to Satisfy The Connectivity Constraints
It should be noted that there are multiple ways to perform layout and routing for this quantum logic circuit. When compiling this quantum logic circuit, QPanda3 may not always use this specific layout and routing scheme, as other layout and routing solutions can still satisfy the target requirements.
V-D Optimization
QPanda3 adopts a strategy that combines local and global optimization. Local optimization primarily focuses on optimizing small regions within the quantum circuit, such as merging adjacent quantum gates and eliminating redundant operations, to reduce the complexity and depth of the circuit. Global optimization, on the other hand, considers the optimization of the entire quantum circuit by adjusting the order of quantum gates, reallocating resources, and other means to further improve the execution efficiency of the circuit.
V-D1 Sabre
Sabre[26] is an advanced quantum circuit compilation algorithm. This algorithm incorporates a decay effect and employs a bidirectional heuristic search based on swaps. It achieves a good trade-off between the depth of the quantum circuit and the number of gates. QPanda3 has further optimized the implementation of Sabre and extensively applies it in the compilation process. Sabre plays a crucial role in optimization, layout, and routing, and through efficient algorithm design, it can handle complex quantum circuits, improving the accuracy and efficiency of compilation. The Sabre algorithm is closely integrated with the characteristics of quantum hardware, enabling the compiled circuit to better adapt to the hardware’s execution environment, thereby ensuring efficient execution of the algorithm.
V-D2 Independent Optimization Steps
This step involves isolating certain optimization strategies as a standalone procedure, intended for pre-layout and post-routing optimization. The optimization strategies concerned encompass merging adjacent quantum gates, eliminating redundant operations, simplifying complex gate sequences, and so forth. These measures aim to reduce the depth and complexity of the circuit, thereby enhancing execution efficiency. The following two simplistic examples illustrate the optimization effects. It is noteworthy that the examples utilize gates such as RZ and CZ, which are not decomposed or transformed during the preprocessing stage.
(1) Example 1 - Merging Adjacent Gates
This circuit presents a scenario that is amenable to optimization, where consecutive RZ gates can be consolidated. This scenario can be further generalized to the optimization of various types of consecutive single-gate operations. In this instance, when QPanda3 compiles the circuit at optimization level 0, no optimization strategies are employed. Consequently, the result retains two consecutive RZ gate operations, consistent with the quantum logic circuit as depicted in the figure. However, when QPanda3 compiles the circuit at optimization levels 1 and 2, it merges these two consecutive RZ gates into a single RZ gate. The optimized circuit necessitates only the time required to execute one RZ gate, whereas the pre-optimized circuit consumes approximately twice the duration.
Figure 9: Merging Adjacent Gates
(2) Example 2 - Eliminating Redundant Gates
This circuit showcases a scenario that can be optimized, where consecutive SWAP gates exist and can be eliminated. This scenario can be extended to the optimization of various types of consecutive two-qubit gate operations. When QPanda3 performs quantum circuit compilation at optimization level 2, adjacent SWAP gates are removed. The elimination of redundant gates effectively enhances the utilization of quantum bits.
Figure 10: Eliminating Redundant Gates
V-D3 Pre-placement Optimization and Post-routing Optimization
Prior to the placement stage, QPanda3 conducts a series of preprocessing optimizations on the quantum circuit. These optimization operations aim to alleviate the burden on subsequent placement and routing processes, thereby improving compilation efficiency. For instance, by merging adjacent quantum gates, the number of gates that need to be considered during the placement stage can be reduced. Additionally, by eliminating ineffective operations, the occupation of quantum hardware resources can be minimized. After the routing stage is completed, QPanda3 further optimizes the circuit to enhance its execution efficiency even more.
VI Device-Based Quantum Program Analysis
Whether a quantum program can be executed efficiently on a specific computing device and yield desirable results determines its suitability for performance-sensitive tasks and its advantage in terms of computational resource costs. Efficient utilization of device resources in quantum programs relies on device-oriented fine-tuning. On one hand, skillful program design based on programming experience can be employed; on the other hand, purposeful adjustments can be made according to the runtime performance of the quantum program on the computing device. This section focuses on the latter, introducing the tools provided by QPanda3 for device-based quantum program analysis.
VI-A Runtime State Information and Quantum Program Analysis
In classical computing, runtime state information of various hardware components can be directly obtained through dedicated counters, registers, and debug-level interfaces during program execution. Due to the unique nature of quantum mechanics, however, observing a quantum processor without interrupting its operation poses challenges. Typically, one cannot read information from physical qubits as one would from bits in a classical register without causing the program to halt. Nevertheless, in the NISQ (Noisy Intermediate-Scale Quantum) era, a necessary condition for quantum computing is the ability to drive the directed evolution of qubits externally. Information such as the type and duration of externally initiated drive operations can clearly be acquired and fed back to the software layer. Additionally, statistical information on certain runtime-inaccessible parameters can be obtained through multiple trials. This quantum device information can be published as product specification parameters on the one hand, and re-collected as needed on the other. In the context of quantum program analysis, the former is suitable for analyzing standard product specifications, while the latter is applicable to analyzing the current state of the device.
Software is an essential component for program analysis. Statistical information about the hardware can be used as input data for direct analysis by the software. Alternatively, runtime-related information from the hardware can be collected within the software itself to analyze the program’s performance on the current hardware during the present time period.
QPanda3 is a modern software tool designed to serve quantum computing. It provides device-oriented quantum program analysis tools. The open interfaces allow users to perform device-specific quantum program analysis based on the specification parameters and runtime statistical information of quantum computing devices. Integrated within the quantum cloud, QPanda3 operates in high synergy with quantum computers to meet the demands of quantum program analysis for various objectives. Currently, QPanda3 offers quantum circuit analysis tools tailored to device performance and quantum program performance analysis tools based on workflow and device information.
VI-B Quantum Circuit Analysis for Device Performance
VI-B1 Overview
QPanda3 extends the quantum computing performance benchmarks to the quantum circuit analysis context. Specifically, QPanda3 incorporates key benchmarks proposed by SupermarQ[35], including Program Communication, Critical-Depth, Entanglement-Ratio, Parallelism, and Liveness. These benchmarks serve to measure the quantum computing performance of quantum circuits on specific devices, providing guidance for users to evaluate and improve quantum programs.
When utilizing the open interfaces provided by QPanda3, users are required to supply the compiled circuit as the object of analysis. The compiled circuit can be obtained through the circuit compilation tools introduced in Section 5.
QPanda3 packages and returns the quantum program analysis results based on the aforementioned benchmarks in the form of Python lists. This approach facilitates users in collecting corresponding result data when analyzing a large number of circuits. Additionally, QPanda3 provides corresponding visualization tools.
Here, an example is provided to illustrate how to use QPanda3 to obtain the performance metrics of a quantum program, including Program Communication, Critical-Depth, Entanglement-Ratio, Parallelism, and Liveness. QPanda3 offers quantum program analysis functionality through the draw_circuit_features interface in the pyqpanda3.profiling package. The following code constructs a simple quantum circuit, which is managed using a QCircuit object. Finally, the draw_circuit_features function is utilized to perform the analysis.
from pyqpanda3.profiling import draw_circuit_features
from pyqpanda3.core import QCircuit, H, CNOT, X
circuit = QCircuit(4)
circuit << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3)
draw_circuit_features(circuit, True)
The corresponding visualization results are shown in Figure 11.
Figure 11: Performance Metrics of A Quantum Program
VI-C Performance Analysis of Quantum Programs Based on Flow and Device Information
Inspired by classical program analysis methods, QPanda3 provides users with a tool for analyzing quantum programs based on flow and device information. This tool assists users in analyzing the utilization rates and execution time percentages of various quantum logic gates within their respective processes in quantum programs, providing a basis for users to evaluate and improve their quantum programs. This tool is implemented based on gprof. Next, we will first demonstrate the usage of relevant interfaces through a code example. Then, centering around this example, we will introduce the differences between this feature of QPanda3 and classical program analysis, as well as the other software components of this functionality.
VI-C1 An Example
QPanda3 provides the functionality for performance analysis of quantum programs based on flow and device information through the draw_circuit_profile interface in the pyqpanda3.profiling package. In the following code, two quantum circuits, cir1 and cir2, are first constructed, and then the draw_circuit_profile function is used to analyze the circuit cir2. QPanda3 abstracts circuits consisting solely of quantum gates using the QCircuit class. Adding one QCircuit object to another represents the addition of a new quantum sub-circuit. The device information used includes the quantum gates supported by the quantum processor employed and the average execution time of these gates on the quantum processor. This information is passed as the second parameter to the draw_circuit_profile function in the form of a Python dictionary.
from pyqpanda3.profiling import draw_circuit_profile
from pyqpanda3.core import QCircuit, H, CNOT, X
cir1 = QCircuit(4)
cir1 << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3) << X(3)
cir2 = QCircuit(4)
cir2 << X(1) << cir1
draw_circuit_profile(cir2, {’CNOT’: 200, ’H’: 60, ’X’: 30}, True)
In this code, the third parameter of the draw_circuit_profile function controls whether visualization is performed. The corresponding visualization result is shown in Figure 12.
Figure 12: Profile of Circuit
VI-C2 Comparison with Classical Program Analysis
Quantum sub-circuits bear similarities to called functions (subroutines) in classical programs. Both represent the totality of program operations within a certain period of time. The former represents all quantum logic gate operations within a period, while the latter represents all classical program operations within a period. Both can also contain other sequences of operations. A quantum sub-circuit can contain other sub-circuits, and a called function (subroutine) can call other functions (subroutines). Therefore, both can be effectively described using directed graphs. Function (subroutine) calls can be accurately described using edges in a directed graph. Similarly, the inclusion relationships between sub-circuits can also be accurately described using edges in a directed graph. From the example above, it can be seen that QPanda3 accurately describes the relationships between sub-circuits and between circuits and quantum gates using a directed graph. In Figure 11, QCircuit_0 corresponds to the first instantiated QCircuit object, cir1, and QCircuit_1 corresponds to the second instantiated QCircuit object, cir2. The arrow from QCircuit_1 to QCircuit_0 indicates that the quantum circuit corresponding to QCircuit_1 (cir2) has the quantum circuit corresponding to QCircuit_0 (cir1) as a direct sub-circuit. The arrow from QCircuit_0 to H indicates that the quantum circuit corresponding to QCircuit_0 (cir1) directly contains the H gate. Obviously, the H gate in the diagram is also a sub-circuit of QCircuit_1 (cir2). For clarity, the term ”direct” is used here to specifically refer to relationships like those between QCircuit_0 and QCircuit_1 (or between QCircuit_0 and H).
During execution, there is a fundamental distinction between quantum sub-circuits and called functions (subroutines) in classical programs. In a classical subroutine, each instruction reutilizes the same hardware computing resources, such as the program instruction pointer register, stack register, and program status word register. The program status word register is employed to facilitate subroutine calls in classical programs. However, to prevent the collapse of the quantum system, quantum computing processors that handle entanglement will map all operations in a quantum circuit onto the quantum processor in a single step. That is, sub-circuits like cir1 are fully expanded when cir2 is executed, rather than being invoked and executed through instruction jumps and instruction stack maintenance as in classical programs.
Flow-based quantum program analysis holds significant importance. In the actual execution of quantum circuits, sub-circuits are unfolded, yet their utilization brings great convenience during the design phase of quantum logic circuits. On the one hand, some sub-circuits are adaptations of classical subroutines. On the other hand, some sub-circuits are sequences of quantum gates designed for specific functions. Through flow-based and device-informed quantum program performance analysis, users can ascertain the resource usage of various quantum sub-circuits and quantum gates during execution, thereby adjusting the implementation schemes of corresponding sub-circuits accordingly.
VI-C3 Output Analysis Results
The analysis results provided by draw_circuit_profile offer the frequency and time consumption percentage of each sub-circuit and quantum gate within the quantum circuit. The percentages in Figure 11 represent the corresponding time consumption percentages, and the ”Nx” notation indicates the frequency value N. When multiple arrows converge on a single node, the respective usage frequencies and time consumption percentages are summarized using a summation approach.
QPanda3 also supports outputting results in the form of gprof-style reports. These outputs provide more specific and detailed analysis data. Figure 13 presents partial results from the aforementioned example code.
Figure 13: Gprof-Style Report
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) Visualization of the circuit
Figure 14: For Circuit Count
VII Experiments
VII-A Performance Experiments of OriginBIS
QPanda3 has specifically designed OriginBIS for quantum program transmission. A typical application scenario involves users designing quantum programs on local devices and executing them using abundant computing resources on quantum clouds. In this scenario, quantum programs exist as memory objects in high-level programming languages on the source device. After transmission, they exist as directly executable machine instructions on the target device. This scenario comprises three stages: firstly, the source device converts the quantum program in the form of a high-level programming language memory object into an intermediate representation of the quantum program; secondly, the intermediate representation is transmitted to the target device using a communication link; and finally, the target device converts the intermediate representation into directly executable machine instructions. It should be noted that the scope of the concept of quantum program transmission may vary depending on specific business requirements. If the business only requires the quantum program on the source device to appear on the target device, then transmitting the data storing the intermediate representation of the quantum program from the source device to the target device completes the quantum program transmission, which only includes the second stage. If the business requires quantum program transmission from a memory object on the source device to executable machine instructions on the target device, then the quantum program transmission should encompass all three stages. For convenience of description, in the experiments of this paper, the quantum program transmission process refers to the second stage, the first stage is termed the Encode stage, and the third stage is termed the Decode stage. The experiments in this paper primarily focus on testing and comparing the performance of OriginBIS across these three stages.
This paper tests the performance of OriginBIS in quantum program transmission and quantum program format conversion through multiple comparative experiments. The experimental results demonstrate that OriginBIS significantly improves the processing efficiency of the corresponding processes.
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 15: For Circuit Depth
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 16: For Qubit Count
VII-A1 Evaluation Metrics
(1) Encode Time and Decode Time
For the Encode and Decode stages, the speed at which the device completes the entire processing operations of the respective stages reflects the processing efficiency of the corresponding processes. Considering that it is difficult to quantify the amount of processing operations performed by the device into specific numerical values, this paper uses the time taken to complete the same conversion task to reflect the impact of the intermediate representation on processing efficiency during the Encode and Decode stages. Specifically, for the same conversion task, the longer the time taken, the lower the processing efficiency; the shorter the time taken, the higher the processing efficiency.
(2) Post-encoding Size
Unlike the Encode and Decode stages, where time is used as a metric, this paper employs the data size of the quantum program after encoding (Post-encoding Size) to measure the impact of the intermediate representation on transmission efficiency during the quantum program transmission stage. The quantum program transmission stage constitutes a classical communication process, where latency is a crucial performance evaluation metric. However, in practical communication environments, latency is influenced by various factors, including the bandwidth and distance of physical communication links and devices, as well as the classical network protocols used for data transmission. Quantum program transmission primarily involves transmitting the intermediate representation of the quantum program as pure data. The data size of the intermediate representation is the only component of the quantum program intermediate representation that is directly related to communication latency, excluding the influences of physical communication links, devices, and classical network protocols. Specifically, for the same quantum program transmission task, a larger Post-encoding Size results in lower transmission efficiency, while a smaller Post-encoding Size leads to higher transmission efficiency.
VII-A2 Experimental Setup
This paper conducts comparative experiments from three aspects: the number of circuits, circuit depth, and the number of qubits. Batch transmission of a large number of circuits is one of the real and essential practical application requirements. Circuit depth is a crucial factor that affects the execution results of quantum circuits. The number of qubits measures the amount of quantum computing resources occupied by quantum circuits. These three metrics serve as important references for users when submitting quantum programs and for devices when processing quantum programs for execution. The experiments in this paper are carried out using the scheme shown in Table I. For the experiment on the number of quantum circuits, the number of quantum circuits is taken as the independent variable, while the circuit depth is controlled at a fixed value of 500 and the number of qubits is controlled at a fixed value of 72. For the experiment on circuit depth, the circuit depth is taken as the independent variable, with the number of circuits controlled at a fixed value of 500 and the number of qubits controlled at a fixed value of 72. For the experiment on the number of qubits, the number of qubits is taken as the independent variable, with the circuit depth controlled at a fixed value of 500 and the number of circuits controlled at a fixed value of 500. It should be noted that the parameters related to circuits are not limited to the number of circuits, circuit depth, and the number of qubits. Moreover, there may be numerical constraints among these parameters. For example, when the number of quantum gates is fixed, there is a certain negative correlation between circuit depth and the number of qubits. To minimize the influence of other parameters and the relationships among them, this paper adopts a randomized technical approach, generating random circuits based on the conventions shown in Table I to test the performance of quantum program intermediate representations in quantum program transmission. Relevant experiments are conducted on OriginBIS, OriginIR, and QASM to compare the quantum program transmission performance of these intermediate representations.
TABLE I: Experimental Setup For Quantum Program Transmission
Experiment Circuit Count Circuit Depth Qubit Count
For Circuit Count Independent Variable 500 72
For Circuit Depth 500 Independent Variable 72
For Qubit Count 500 500 Independent Variable
VII-A3 Experimental Results and Analysis
(1) For Circuit Count
The experimental results for the number of circuits are shown in Figure 14. The grouped columns chart in subplots (a), (b), and (c) demonstrates that OriginBIS is significantly smaller than OriginIR and QASM in terms of Encode Time, Decode Time, and Post-encoding Size. This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in the efficiency of quantum program transmission. Through simple numerical estimation, it can be found that for Encode Time, OriginIR is approximately 17 times that of OriginBIS, and QASM is approximately 80 times that of OriginBIS; for Decode Time, OriginIR is approximately 20 times that of OriginBIS, and QASM is approximately 40 times that of OriginBIS; for Post-encoding Size, OriginIR is approximately 3.5 times that of OriginBIS, and QASM is approximately 6.5 times that of OriginBIS. The point-line charts in subplots (d), (e), and (f) show that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth in the number of circuits.
(2) For Circuit Depth
The experimental results regarding circuit depth are presented in Figure 15. The grouped columns chart in subplots (a), (b), and (c) demonstrate that OriginBIS exhibits significantly lower values in terms of Encode Time, Decode Time, and Post-encoding Size compared to both OriginIR and QASM. This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission. Upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 18 times that of OriginBIS, while QASM is about 88 times; for Decode Time, OriginIR is roughly 21 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times. The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth of circuit depth.
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 17: For Gate Count
(3) For Qubit Count
The experimental results concerning the number of qubits are illustrated in Figure 16. The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM. Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM. This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission. Excluding the 90-qubit data in subplot (b) and the 40-bit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times. The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation. However, Post-encoding Time consistently increases with the increase in the number of qubits.
(4) Gate Count
During the experiments examining circuit count, circuit depth, and qubit count, it was observed that the number of quantum gates varies with changes in these parameters. Figure 18 illustrates the trend in the number of quantum gates, demonstrating a strong positive linear correlation with the value of the third parameter when the values of any two of these three quantities are held constant.
The number of gates is also an important metric for evaluating quantum programs. By combining the gate count data from Figure 18 with the corresponding Encode Time, Decode Time, and Post-encoding Size data, the results presented in Figure 17 were obtained. Subplots (a), (b), and (c) indicate that the Encode Time, Decode Time, and Post-encoding Size for OriginBIS are significantly lower than those for OriginIR and QASM. Subplots (d), (e), and (f) reveal that there is a trend of increasing Encode Time and Decode Time with the growth in the number of quantum gates, albeit with noticeable data fluctuations. However, Post-encoding Time increases with the increase in the number of quantum gates, exhibiting a strong linear characteristic.
The experimental results concerning the number of qubits are illustrated in Figure 16. The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM. Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM. This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission. Excluding the 90-qubit data in subplot (b) and the 40-qubit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times. The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation. However, Post-encoding Time consistently increases with the increase in the number of qubits.
Figure 18: Gate Count for Circuit Count, Circuit Depth and QBit Count
Figure 19: Experimental Results About Circuit Compilation
VII-B Compilation Efficiency Comparison Experiment
VII-B1 Experiment Setup
In this paper, a comparative experiment related to compilation efficiency was conducted on multiple mainstream quantum programming frameworks based on Benchpress. Benchpress is an evaluation suite for multi-quantum computing software development kits, containing thousands of test cases. It allows for the uniform testing of multiple quantum software packages across various performance and functional indicators, with the results reflecting the cost of processing quantum circuits on quantum computing devices using different software packages. All relevant experiments for each software package mentioned in the Benchpress literature were replicated in this study, with the addition of tests for QPanda3. All software packages tested are listed in Table II. It should be noted that a timeout limit of 180 seconds was imposed for executing each test case.
The testing of these software packages was conducted in the same software and hardware environments. The environmental parameters used in our experiments are presented in Table III.
TABLE II: SDKs
SDK Version
amazon-braket-sdk(braket) 1.91.0
bqskit 1.2.0
cirq 1.4.1
pytket(tket) 2.0.1
qiskit 1.4.2
qiskit_ibm_transpiler 0.10.3
pystaq 3.5
pyqpanda3 0.2.0
TABLE III: Software and Hardware Environments
Type Info
CPU Intel(R) Xeon(R) Platinum 8336C CPU @ 2.30GHz
Memory 2.0TB
Operating System Ubuntu 20.04.5 LTS
Python 3.11.11
VII-B2 Evaluation Metrics
This paper adopts the evaluation metrics introduced by Benchpress. They are Standard Pytest Output Type, circuit construction time, manipulate time, and circuit compilation time.
VII-B3 Experimental Results And Analysis
TABLE IV: Experimental Results Evaluated Based on The Criterion of Standard Pytest Output Type
SDK Passed Failed XFailed Skipped Total
braket 7 2 0 1057 1066
cirq 10 2 0 1054 1066
qiskit 1044 0 0 22 1066
QPanda3 1016 28 0 22 1066
In our experiments, the results of the test_status_counts are presented in Table IV. It was observed that many SDKs failed to satisfy the requirements of numerous test cases in Benchpress. Specifically, bqskit was only capable of performing tests related to circuit construction. pytket and pystaq exhibited compatibility issues with the environments of other SDKs, making it impossible to conduct fair testing and comparison using a unified test suite and environment. Furthermore, qiskit_ibm_transpiler demonstrated strong dependency on the IBM computing platform, rendering it untestable under our experimental environment. As evident from Table IV, both braket and cirq skipped a significant number of test cases. According to the Benchpress literature, qiskit skipped 22 test cases and passed 1,044 test cases. Tabel IV indicates that our experiments actually tested multiple SDKs using 1,066 test cases. The performance of qiskit in our experiments was consistent with the reports in the Benchpress literature. QPanda successfully passed $95.3\%$ of all test cases, with the number of skipped cases equal to that of qiskit. This table showcases the excellent versatility of QPanda3 within the unified test suite, Benchpress.
The performance of each SDK in circuit construction is illustrated in Figure 20. It can be observed that QPanda3, qiskit, and cirq passed all seven test cases. For the five test cases, namely test_multi_control_circuit, test_bigint_qasm2_import, test_param_circSU2_100_build, test_clifford_build, and test_QV100_build, QPanda3 demonstrated the shortest execution time. For the test case test_param_circSU2_100_bind, QPanda3’s execution time was almost on par with qiskit, ranking as the second fastest. In the test case test_QV100_qasm2_import, although QPanda3 lagged behind qiskit, its execution time was significantly less than that of cirq and bqskit. Despite QASM not being the native intermediate representation for quantum programs in QPanda3, it still exhibited a circuit construction speed close to that of qiskit in these two test cases. This suggests that the QASM conversion tool in QPanda3 may be highly efficient, or that QPanda3’s circuit construction efficiency is so high that it can offset any inefficiencies in QASM conversion. This figure highlights the advantages of QPanda3 in circuit construction.
Figure 20: Experimental Results About Circuit Construction(shorter is better)
Figure 21: Experimental Results About Circuit Manipulation(shorter is better)
The Benchpress literature highlights that Manipulate is a crucial metric for evaluating SDKs. As shown in Figure 21, only QPanda3 and qiskit passed all four test cases. QPanda3 demonstrated the shortest execution time in three tests: test_multi_control_decompose, test_QV100_basis_change, and test_random_clifford_decompose. For the test_DTC100_twirling, QPanda3’s execution time was only slightly longer than that of qiskit, with a minimal difference between the two. This figure indicates that QPanda3 has a clear advantage in the Manipulate metric.
In our experiments, only QPanda3 and qiskit passed the majority of the test cases related to compilation. We have plotted the corresponding experimental results into multiple scatter plots, as shown in Figure 19, which illustrate the differences between QPanda3 and qiskit in terms of 2Q Gate Count, 2Q Gate Depth, and compilation time for various topological structures. In each plot, the blue dashed line represents the point where the values for QPanda3 and qiskit are equal for the corresponding metric. Points above the blue dashed line indicate that the corresponding value for QPanda3 is less than that for qiskit. As observed from Figure 19, for both 2Q Gate Count and 2Q Gate Depth, the majority of the scatter points are densely distributed along or very close to the blue dashed line, suggesting that the performance of QPanda3 and qiskit is essentially the same. However, regarding compilation time, all scatter points in the four subplots of Figure 19 are located above the blue dashed line, indicating that QPanda3 exhibits higher compilation efficiency than qiskit. This characteristic is not limited by topological structure and is effective for a large number of different circuits. Additionally, it can be observed that only a very few test cases result in compilation times exceeding 10 seconds when using QPanda3. These observations collectively demonstrate the high compilation efficiency of QPanda3.
VIII Conclusion
This paper focuses on efficient compilation and execution in quantum computing, introducing OriginIR and OriginBIS as intermediate representations to enhance programmability, transmission efficiency, and execution performance. Experimental results demonstrate that OriginBIS significantly outperforms OpenQASM 2.0 in encoding and decoding speed as well as information capacity. Additionally, QPanda3 surpasses Qiskit in quantum circuit construction, operation execution, and compilation speed, exhibiting exceptional performance improvements, particularly in large-scale quantum circuit compilation tasks. These advancements lay a solid foundation for the engineering applications of future quantum computing.
IX Future Work
The quantum programming framework QPanda3 is poised to enhance its capabilities for future ”quantum-HPC-AI” integrated computing by focusing on several key advancements. Building on its existing strength of compiling quantum circuits 14.97 times faster than Qiskit in Benchpress tests, QPanda3 will develop hybrid task schedulers to enable seamless coordination between quantum computations and classical supercomputing resources. It aims to integrate AI-driven optimization tools, such as AutoML modules, to automatically reduce quantum gate counts and circuit depths while mitigating noise impacts. The framework will expand its hybrid programming interfaces to support classical HPC technologies like MPI and CUDA, facilitating efficient quantum-classical algorithm interoperability. Additionally, QPanda3 plans to implement a unified resource management platform for dynamic allocation of quantum hardware, supercomputing clusters, and AI accelerators. To strengthen ecosystem integration, it will promote standardization through open-source collaboration, enabling deeper compatibility with classical AI libraries (e.g., TensorFlow) while maintaining its low-latency compilation advantages. These upgrades will position QPanda3 as a core enabler for cross-paradigm applications in finance, drug discovery, and materials science.